# Continual Learning for Large Language Models: A Survey

### 02 Feb 2024 v1

@@ Abstract
Large language models (LLMs) are not amenable to frequent re-training, due to high training costs arising from their massive scale. However, updates are necessary to endow LLMs with new skills and keep them up-to-date with rapidly evolving human knowledge. This paper surveys recent works on continual learning for LLMs. Due to the unique nature of LLMs, we catalog continue learning techniques in a novel multi-staged categorisation scheme, involving continual pretraining, instruction tuning, and alignment. We contrast continual learning for LLMs with simpler adaptation methods used in smaller models, as well as with other other enhancement strategies like retrieval-augmented generation and model editing. Moreover, informed by a discussion benchmarks and evaluation, we identify a number of challenges and future work directions for this crucial task.
@@ 2.3 Continual Learning for LLMs
Besides categorizing methods based on training stages, we also provide an alternative categorization based on the information updated during continual learning. In Table 2, we list some representative information is updated for LLMs, e.g., facts, domains, tasks, values, and preferences. Based on the training objectives of LLMs, this information can be updated in different stages of LLM continual learning. The taxonomy in Figure 3 shows our categorization scheme and recent representative work in each category.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=7pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=11em,font=,, where level=3text width=11em,font=,, where level=4text width=10em,font=,, [ Continual Learning for Large Language Models,fill=hidden-yellow!70,ver [ Continual Pre-training (§3),fill=hidden-red!70 [ Update Fact (§3),fill=hidden-blue!70 [ Jang et al. (2022b), Sun et al. (2020), Jang et al. (2022a), leaf, text width=48.5em,fill=hidden-blue!70 ] ] [ Update Domain (§3),fill=hidden-blue!70 [ Domain-Incremental
@@
Continual Pre-training,fill=hidden-blue!70 [ Yang et al. (2024), Xie et al. (2023), Ma et al. (2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Update Language (§3),fill=hidden-blue!70 [ Natural Language,fill=hidden-blue!70 [ Castellucci et al. (2021), Gogoulou et al. (2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Programming Language,fill=hidden-blue!70 [ Yadav et al. (2023), Zan et al. (2022), leaf, text width=38em,fill=hidden-blue!70 ] ] ] ] [ Continual Instruction Tuning (§4),fill=hidden-red!70 [ Update Domain (§4),fill=hidden-blue!70 [ Wang et al. (2023b), Song et al. (2023),Cheng et al. (2023a), Cheng et al. (2023b), Zhang et al. (2023c) , leaf, text width=48.5em,fill=hidden-blue!70 ] ] [ Update Task (§4),fill=hidden-blue!70 [ Razdaibiedina et al. (2023), Anonymous (2024b), Mok et al. (2023), Jang et al. (2023), Wang et al. (2023a) , leaf, text width=48.5em,fill=hidden-blue!70 ] ] [ Update Skill (§4),fill=hidden-blue!70 [ Hao et al. (2023), Kong et al. (2023), Qin et al. (2023a), Jin et al. (2023), , leaf, text width=48.5em,fill=hidden-blue!70 ] ] ] [ Continual Alignment (§5),fill=hidden-red!70 [ Update Values (§5),fill=hidden-blue!70 [ Zhang et al. (2023a), Anonymous (2024a) , leaf, text width=48.5em,fill=hidden-blue!70 ] ] [ Update Preference (§5),fill=hidden-blue!70 [ Suhr and Artzi (2023) , leaf, text width=48.5em,fill=hidden-blue!70 ] ] ] ]

### 07 Feb 2024 v2

@@ Abstract
Large language models (LLMs) are not amenable to frequent re-training, due to high training costs arising from their massive scale. However, updates are necessary to endow LLMs with new skills and keep them up-to-date with rapidly evolving human knowledge. This paper surveys recent works on continual learning for LLMs. Due to the unique nature of LLMs, we catalog continue learning techniques in a novel multi-staged categorization scheme, involving continual pretraining, instruction tuning, and alignment. We contrast continual learning for LLMs with simpler adaptation methods used in smaller models, as well as with other enhancement strategies like retrieval-augmented generation and model editing. Moreover, informed by a discussion of benchmarks and evaluation, we identify a number of challenges and future work directions for this crucial task.
@@ 2.3 Continual Learning for LLMs
Besides categorizing methods based on training stages, we also provide an alternative categorization based on the information updated during continual learning. In Table 2, we list some representative information that is updated for LLMs, e.g., facts, domains, tasks, values, and preferences. Based on the training objectives of LLMs, this information can be updated in different stages of LLM continual learning. The taxonomy in Figure 3 shows our categorization scheme and recent representative work in each category.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=7pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=11em,font=,, where level=3text width=11em,font=,, where level=4text width=10em,font=,, [ Continual Learning for Large Language Models,fill=hidden-yellow!70,ver [ Continual Pre-training (§3),fill=hidden-red!70 [ Update Fact (§3.1),fill=hidden-blue!70 [ Jang et al. (2022b), Sun et al. (2020), Jang et al. (2022a), leaf, text width=48.5em,fill=hidden-blue!70 ] ] [ Update Domain (§3.2),fill=hidden-blue!70 [ Domain-Incremental
@@
Continual Pre-training,fill=hidden-blue!70 [ Yang et al. (2024), Xie et al. (2023), Ma et al. (2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] ] [ Update Language (§3.3),fill=hidden-blue!70 [ Natural Language,fill=hidden-blue!70 [ Castellucci et al. (2021), Gogoulou et al. (2023) , leaf, text width=38em,fill=hidden-blue!70 ] ] [ Programming Language,fill=hidden-blue!70 [ Yadav et al. (2023), Zan et al. (2022), leaf, text width=38em,fill=hidden-blue!70 ] ] ] ] [ Continual Instruction Tuning (§4),fill=hidden-red!70 [ Update Task (§4.1),fill=hidden-blue!70 [ Razdaibiedina et al. (2023), Anonymous (2024b), Mok et al. (2023), Jang et al. (2023), Wang et al. (2023a) , leaf, text width=48.5em,fill=hidden-blue!70 ] ] [ Update Domain (§4.2),fill=hidden-blue!70 [ Wang et al. (2023b), Song et al. (2023),Cheng et al. (2023a), Cheng et al. (2023b), Zhang et al. (2023c) , leaf, text width=48.5em,fill=hidden-blue!70 ] ] [ Update Skill (§4.3),fill=hidden-blue!70 [ Hao et al. (2023), Kong et al. (2023), Qin et al. (2023a), Jin et al. (2023), , leaf, text width=48.5em,fill=hidden-blue!70 ] ] ] [ Continual Alignment (§5),fill=hidden-red!70 [ Update Values (§5.1),fill=hidden-blue!70 [ Zhang et al. (2023a), Anonymous (2024a) , leaf, text width=48.5em,fill=hidden-blue!70 ] ] [ Update Preference (§5.2),fill=hidden-blue!70 [ Suhr and Artzi (2023) , leaf, text width=48.5em,fill=hidden-blue!70 ] ] ] ]